1. The revenue R in dollars, derived from selling x surf boards is given by
R = -2 x^2 + 4800 x +1000

How many items should be sold to maximize the revenue?
What is the revenue?

Answer by nerdybill(7384)

1. The revenue R in dollars, derived from selling x surf boards is given by
R = -2 x^2 + 4800 x +1000
How many items should be sold to maximize the revenue?
By inspecting the coefficient associated with the x^2 term, we know that it is a parabola that opens downward. Therefore, the vertex represents the max.
The 'x' value of the vertex is:
x = -b/(2a)
x = -4800/(2*(-2))
x = -4800/(-4)
x = 1200
.
What is the revenue?
R = -2 x^2 + 4800 x +1000
R = -2(1200)^2 + 4800(1200) +1000
R = $2,881,000